# Relaxed Equivariance via Multitask Learning

### v2

@@ 5 Related Work
Equivariant ML Models. In the vision domain, group convolutions have proven to be a powerful tool for handling rotation equivariance for images and enhanced model generalization (Cohen & Welling, 2016; Cohen et al., 2019; Weiler & Cesa, 2019; Qiao et al., 2023). Similarly, the development of equivariant architectures with respect to roto-translations for geometric data has been an active area of research (Chen et al., 2021a; Satorras et al., 2021; Han et al., 2022; Xu et al., 2024). Techniques that use spherical harmonics and irreducible representations have shown a large success in modeling graph-structured data, such as SE(3)-Transformers (Fuchs et al., 2020), Tensor Field Networks (Thomas et al., 2018), and DimeNet (Gasteiger et al., 2020). More recently, Brehmer et al. (2023) introduced an E(3) equivariant Transformer that employs geometric algebra for processing 3D point clouds.
@@
Data Augmentation and Unconstrained Models. Alternatively, integrating transformations through data augmentation is a widely used strategy across multiple vision tasks, enhancing performance in image classification (Perez & Wang, 2017; Inoue, 2018; Rahat et al., 2024), object detection (Zoph et al., 2020; Wang et al., 2019; Kisantal et al., 2019), and segmentation (Negassi et al., 2022; Chen et al., 2021b; Yu et al., 2023). For geometric data, Hu et al. (2021) has adapted a Graph Neural Network architecture with data augmentation to process 3D molecular structures. In parallel, Dosovitskiy et al. (2021) introduced that Vision Transformers (ViTs) with a large amount of training data can achieve comparable performance with Convolutional Neural Networks (CNNs), obviating the need for explicit translation equivariance within the architecture. Recently, this has shown to be effective for handling geometric data (Wang et al., 2024; Abramson et al., 2024).
@@
Learning Symmetries and Approximate Equivariance. Several studies have shown that the layers of CNN architectures can be approximated for a soft constraint (Wang et al., 2022; van der Ouderaa et al., 2022; Romero & Lohit, 2022; Veefkind & Cesa, 2024; Wu et al., 2024; McNeela, 2023). Conversely, van der Ouderaa et al. (2023) extends the Bayesian model selection approach to learning symmetries in image datasets. Yeh et al. (2022) introduced a parameter-sharing scheme to achieve permutations and shifts equivariances in Gaussian distributions. Recent works have relaxed the hard constrained models to a soft constraint by adding unconstrained layers in the architecture design (Finzi et al., 2021a; Pertigkiozoglou et al., 2024), canonicalization network (Lawrence et al., 2024). , or explicit relaxation (Kaba & Ravanbakhsh, 2023). Additionally, Lin et al. (2019) modified the loss of CNN for segmentation task. Shakerinava et al. (2022) introduced a method to learn equivariant representation using the group invariants, while Bhardwaj et al. (2023) defined a regularizer that injects the equivariance in the latent space of the network by explicitly modeling transformations with additional learnable maps. In contrast, several works have started from pre-trained models (Basu et al., 2023; Kim et al., 2023b). Furthermore, the EGNN framework (Satorras et al., 2021) has been modified using an invariant function (Zheng et al., 2024) or adversarial training procedure (Yang et al., 2023). However, in our work, we start from completely unconstrained models, without imposing any equivariance constraints on the space of functions within the architecture. Moreover, we didn’t assume a specific class of models or introduce additional parameters, which increases the applicability of our method to various domains and makes it computationally efficient.
@@ 6 Experiments and Discussion
Figure 2: N-body dynamical system. Each row represents a different evaluation scenario: The top row shows in-distribution performance, the middle row displays out-of-distribution performance, and the bottom row illustrates equivariance error. The columns correspond to different architectures/ model conditions (from left to right): The first column shows the Transformer trained with REMUL (gradual penalty), the second column with a constant penalty, and the third column presents the baselines (equivariant models, standard Transformer, and data augmentation). The equivariance metric included in this Figure is defined in Equation 12, we report the same plots for the metric defined in Equation 13 in Appendix B.1 (Figure 6), which has a similar behavior. Transformer architecture with high $\beta$ reduces the equivariance error and improves the performance. SE(3)-Transformer and GATr have a small equivariance error below the range of the plots ($2.8e^{-10}$ and $1.13e^{-15}$ respectively).
We evaluate our method on different tasks for geometric data: N-body dynamical system (Section 6.1), motion capture (Section 6.2), and molecular dynamics (Section 6.3). For unconstrained models, we apply REMUL to Transformers and Graph Neural Networks. We then compare against their equivariant baselines: SE(3)-Transformer (Fuchs et al., 2020), Geometric Algebra Transformer (Brehmer et al., 2023), and Equivariant Graph Neural Networks (Satorras et al., 2021) as well as unconstrained models with data augmentation. We consider learning the rotation group $SO(3)$ for REMUL and data augmentation and we subtract the center of mass for translation. We use the equivariance metric defined in Equation 12 to analyze our results. We also conduct a comparative analysis for the computational requirements of unconstrained models and equivariant models in Section 6.4. Lastly, we discuss the loss surfaces in Section 6.5. Implementation details and additional experiments can be found in Appendix A & Appendix B.
@@ 6.1 N-Body Dynamical System
To conduct ablation studies of our method, we utilized the dynamical system problem described by Brehmer et al. (2023). The task involves predicting the positions of particles after $100$ Euler time steps of Newton’s motion equation, given initial positions, masses, and velocities. This problem is inherently equivariant under rotation and translation groups, implying that any rotation/translation of the initial states should rotate/translate the final states of the particles by the same amount. We conduct comparisons between Transformer trained with REMUL against two equivariant architectures: SE(3)-Transformer and Geometric Algebra Transformer (GATr). We use the same Transformer version and hyperparameters specified by Brehmer et al. (2023). Additional implementation details, including in-distribution and out-of-distribution settings, are provided in Appendix A.1. Our results are presented in Figure 2.
Figure 3: Motion Capture dataset: Transformer trained with REMUL. Two figures on the left: Performance (MSE) and equivariance error for walking task (Subject #$35$), respectively. Two figures on the right: Performance (MSE) and equivariance error for running task (Subject #$9$), respectively. We use the equivariance metric described in Equation 12 and include the same plots for the second metric (Equation 13) in Appendix B.2 (Figure 8). We show a trade-off between model performance and equivariance error, where high penalty $\beta$ gives less equivariance error (more equivariant model) but the best performance comes at an intermediate level of equivariance for both tasks.
Table 1: Performance on Motion Capture dataset: MSE (×$\times 10^{-2}$). REMUL procedure and data augmentation were applied to standard Transformer & MLP. First, Second (highlighted). REMUL comes the best in both tasks.
SE(3)-Transformer GATr Transformer Data Augmentation REMUL
Walking (Subject #$35$) $10.85$±$\pm 1.3$ $10.06$±$\pm 1.3$ $\bf 5.21$±$\pm 0.08$ $5.3$±$\pm 0.18$ $\bf 4.95$±$\pm 0.1$
Running (Subject #$9$) $42.13$±$\pm 3.4$ $32.38$±$\pm 3.9$ $\bf 20.78$±$\pm 1.5$ $29.83$±$\pm 1.4$ $\bf 18.5$±$\pm 0.7$
EMLP RPP PER MLP Data Augmentation REMUL
Walking (Subject #$35$) $7.01$±$\pm 0.46$ $6.99$±$\pm 0.21$ $7.48$±$\pm 0.39$ $6.80$±$\pm 0.18$ $\bf 6.37$±$\pm 0.04$ $\bf 6.04$±$\pm 0.09$
Running (Subject #$9$) $57.38$±$\pm 8.39$ $34.18$±$\pm 2.00$ $\bf 33.03$±$\pm 0.37$ $39.56$±$\pm 2.25$ $40.23$±$\pm 0.94$ $\bf 32.57$±$\pm 1.47$
@@ 6.2 Motion Capture
We further illustrate a comparison on a real-world task, the Motion Capture dataset from CMU (2003). This dataset features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$). We adhered to the standard experimental setup found in the literature (Han et al., 2022; Huang et al., 2022; Xu et al., 2024), employing a train/validation/test split of $200$/$600$/$600$ for Walking and $200$/$240$/$240$ for Running. Additional details can be found in Appendix A.2.
@@
We apply our training procedure REMUL to the Transformer architecture and compare it with SE(3)-Transformer, Geometric Algebra Transformer (GATr), standard Transformer, and Transformer trained with data augmentation. We also compare with Equivariant MLP (Finzi et al., 2021b), as well as two approximate equivariance architectures: Residual Pathway Priors (RPP) (Finzi et al., 2021a), and Projection-Based Equivariance Regularizer (PER) (Kim et al., 2023a). As these architectures are designed specifically on MLP and linear layers, we apply our method to a standard MLP with a similar number of parameters. Our results are presented in Table 1. For REMUL, we provide plots on how the performance and equivariance error change w.r.t. the penalty parameter $\beta$ in Figure 3.
Figure 4: MD17 dataset: GNN trained with REMUL. The first row represents model performance (MSE), and the second row shows equivariance error. Columns from left to right show Aspirin, Ethanol, Malonaldehyde, and Uracil, respectively. The equivariance metric shown in this Figure is defined in Equation 12; we include the same plots for the second metric (Equation 13) in Appendix B.3 (Figure 10). The equivariance error decreases on all molecules with a higher value of $\beta$. In contrast, the required equivariance for best model performance varies for each molecule.
@@
Table 2: Performance on MD17 dataset: MSE (×$\times 10^{-2}$). REMUL procedure and data augmentation were applied to GNN. First, Second (highlighted). REMUL comes the best on six molecules and the second on two molecules.
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic Toluene Uracil
EGNN $14.41$±$\pm 0.15$ $62.40$±$\pm 0.53$ $4.64$±$\pm 0.01$ $\bf 13.64$±$\pm 0.01$ $\bf 0.47$±$\pm 0.02$ $\bf 1.02$±$\pm 0.02$ $11.78$±$\pm 0.07$ $\bf 0.64$±$\pm 0.01$
GNN $\bf 9.26$±$\pm 0.40$ $\bf 26.13$±$\pm 0.11$ $\bf 4.26$±$\pm 0.03$ $18.45$±$\pm 0.54$ $\bf 0.54$±$\pm 0.001$ $\bf 1.02$±$\pm 0.02$ $\bf 9.93$±$\pm 0.82$ $0.70$±$\pm 0.001$
Data Augmentation $\bf 13.7$±$\pm 0.04$ $110.93$±$\pm 5.3$ $5.74$±$\pm 0.02$ $\bf 13.65$±$\pm 0.02$ $0.69$±$\pm 0.001$ $1.33$±$\pm 0.04$ $19.14$±$\pm 0.001$ $0.73$±$\pm 0.002$
REMUL $\bf 9.28$±$\pm 0.40$ $\bf 25.95$±$\pm 0.18$ $\bf 4.02$±$\pm 0.16$ $\bf 13.59$±$\pm 0.03$ $\bf 0.54$±$\pm 0.001$ $\bf 0.99$±$\pm 0.001$ $\bf 9.38$±$\pm 0.20$ $\bf 0.67$±$\pm 0.001$
@@
Table 1 indicates that when processing 3D positions related to human motions, both SE(3)-Transformer and GATr perform worse than the standard Transformer. This outcome is noteworthy because human motion inherently lacks symmetry along the vertical or gravity axis. Consequently, the assumption of equivariance across all axes may not be beneficial or even detrimental. In contrast, a standard Transformer trained with REMUL has the best performance in both tasks.
Following Figure 3, there is a noticeable trade-off in model performance with different values of penalty parameter $\beta$. Best performance is observed at an intermediate level of equivariance, where the model balances between being too rigid (fully equivariant) and too flexible (non-equivariant). This finding underscores the importance of carefully considering the specific characteristics of the data and the task when designing equivariant architectures.
@@ 6.3 Molecular Dynamics
We also present a comparative analysis between constrained equivariant models and unconstrained models focusing on molecular dynamics, specifically predicting 3D molecule structures. We utilize the MD17 dataset (Chmiela et al., 2017), which comprises trajectories of eight small molecules. We use the same dataset split in Huang et al. (2022); Xu et al. (2024), allocating $500$ samples for train, $2000$ for validation, and $2000$ for test. For this task, we selected the Equivariant Graph Neural Network (EGNN) architecture and its non-equivariant GNN counterpart, as presented in Satorras et al. (2021). We then apply REMUL procedure as well as data augmentation to the GNN architecture. Both architectures have the same hyperparameters. More information is indicated in Appendix A.3.
Our results are provided in Table 2. We illustrate how the performance and equivariance error of a GNN trained with REMUL vary across different molecules as a function of $\beta$, as shown in Figure 4 & Figure 9.
Figure 5: Computational time for Geometric Algebra Transformer (GATr) and Transformer architectures. Plots from left to right: Combined forward pass, backward pass, and inference time, respectively. GATr has the highest time in all scenarios. At inference, the Transformer trained with equivariance loss and data augmentation matches the standard Transformer.
@@
From the results presented in Table 2, GNN trained with REMUL outperforms EGNN in six out of eight molecules. Interestingly, a standard GNN, without data augmentation or REMUL, surpasses the performance of EGNN for two molecules: Aspirin and Toluene. In Figure 4 & Figure 9, we observe that the optimal performance of each molecule is attained at different values of the penalty parameter $\beta$. For instance, Malonaldehyde exhibits a direct correlation between model performance and equivariance, where a higher $\beta$ yields better performance. Conversely, for most other molecules, there appears to be a pronounced trade-off where the best performance is achieved at a lower value of $\beta$. This is particularly evident with molecules like Aspirin, where a standard GNN architecture outperforms EGNN. We also plot the 3D structures of the eight molecules in Figure 11. Molecules such as Malonaldehyde, characterized by their symmetric components, might be ideally suited for equivariant design. However, this advantage does not apply to all molecules. Aspirin on the other side, might have an asymmetric structure and exhibit a range of interactions and dynamic states that equivariant models might simplify. Consequently, for such molecules, less equivariant models could potentially offer more accurate predictions.

### v3

@@ 5 Related Work
Equivariant ML Models. In the vision domain, group convolutions have proven to be a powerful tool for handling rotation equivariance for images and enhanced model generalization [56, 57, 58, 59]. Similarly, the development of equivariant architectures with respect to roto-translations for geometric data has been an active area of research [60, 5, 3, 4]. Techniques that use spherical harmonics and irreducible representations have shown a large success in modeling graph-structured data, such as SE(3)-Transformers [15], Tensor Field Networks [14], and DimeNet [61]. More recently, [45] introduced an E(3) equivariant Transformer that employs geometric algebra for processing 3D point clouds.
@@
Data Augmentation and Unconstrained Models. Alternatively, integrating transformations through data augmentation is a widely used strategy across multiple vision tasks, enhancing performance in image classification [62, 63, 64], object detection [65, 66, 67], and segmentation [68, 69, 70]. For geometric data, [71] has adapted a Graph Neural Network architecture with data augmentation to process 3D molecular structures. In parallel, [72] introduced that Vision Transformers (ViTs) with a large amount of training data can achieve comparable performance with Convolutional Neural Networks (CNNs), obviating the need for explicit translation equivariance within the architecture. Recently, this has shown to be effective for handling geometric data [37, 38].
@@
Learning Symmetries and Approximate Equivariance. Several studies have shown that the layers of CNN architectures can be approximated for a soft constraint [46, 73, 74, 75, 76, 77]. Conversely, [78] extends the Bayesian model selection approach to learning symmetries in image datasets. [79] introduced a parameter-sharing scheme to achieve permutations and shifts equivariances in Gaussian distributions. Recent works have relaxed the hard constrained models to a soft constraint by adding unconstrained layers in the architecture design [80, 81], canonicalization network [82], or explicit relaxation [83]. Additionally, [84] modified the loss of CNN for segmentation task. [85] introduced a method to learn equivariant representation using the group invariants, while [86] defined a regularizer that injects the equivariance in the latent space of the network by explicitly modeling transformations with additional learnable maps. In contrast, several works have started from pre-trained models [87, 88]. Furthermore, the EGNN framework [5] has been modified using an invariant function [89] or adversarial training procedure [90]. However, in our work, we start from completely unconstrained models, without imposing any equivariance constraints on the space of functions within the architecture. Moreover, we did not assume a specific class of models or introduce additional parameters, which increases the applicability of our method to various domains and makes it computationally efficient.
@@ 6 Experiments and Discussion
$0.01$
$0.1$
$1.0$
$10.0$
$100.0$
$0.01$
$0.1$
$1.0$
$10.0$
$100.0$
(a) REMUL: Gradual penalty
$0.01$
$0.1$
$1.0$
$10.0$
$100.0$
(b) REMUL: Constant penalty
(c) Baselines
Figure 1: N-body dynamical system. Each row represents a different evaluation scenario. Top: in-distribution performance, Middle: out-of-distribution performance, Bottom: equivariance error. The columns correspond to different architectures/ model conditions. (a) Transformer trained with REMUL (gradual penalty), (b) Transformer trained with a constant penalty, (c) Baselines (equivariant models, standard Transformer, and data augmentation). We conclude that Transformer architecture with high $\beta$ reduces the equivariance error and improves the performance.
We evaluate our method on different tasks for geometric data: N-body dynamical system (Section 6.1), motion capture (Section 6.2), and molecular dynamics (Section 6.3). For unconstrained models, we apply REMUL to Transformers and Graph Neural Networks. We then compare against their equivariant counterparts: SE(3)-Transformer [15], Geometric Algebra Transformer [45], and Equivariant Graph Neural Networks [5] as well as unconstrained models with data augmentation. We consider learning the rotation group $SO(3)$ for REMUL and data augmentation and we subtract the center of mass for translation. We use the equivariance metric defined in Eq. 8 to analyze our results, and include the second metric in Appendix D. We also conduct a comparative analysis for the computational requirements of unconstrained models and equivariant models in Section 6.4. Lastly, we discuss the loss surfaces in Appendix A. Implementation details and additional experiments can be found in Appendix C & Appendix D.
(a) MSE: Walking
(b) Equiv. error: Walking
(c) MSE: Running
(d) Equiv. error: Running
Figure 2: Motion Capture dataset: Transformer trained with REMUL. We show a trade-off between model performance and equiv. error, where high penalty $\beta$ gives less equiv. error (more equivariant model) but the best performance comes at an intermediate level of equivariance for both tasks.
@@ 6.1 N-Body Dynamical System
Table 1: N-body dynamical system: MSE ($\times 10^{-3}$). First, Second.
[…]
To conduct ablation studies of our method, we utilized the dynamical system problem described by [45]. The task involves predicting the positions of particles after $100$ Euler time steps of Newton’s motion equation, given initial positions, masses, and velocities. This problem is equivariant under rotation and translation groups, implying that any rotation/translation of the initial states should rotate/translate the final states of the particles by the same amount. We conduct comparisons between Transformer trained with REMUL against two equivariant architectures: SE(3)-Transformer and Geometric Algebra Transformer (GATr). We use the same Transformer version and hyperparameters specified by [45]. Implementation details, including in-distribution and out-of-distribution settings, in Appendix C.2. Our results are presented in Figure 1 and Table 1.
@@ 6.2 Motion Capture
We further illustrate a comparison on a real-world task, the Motion Capture dataset from [91]. This dataset features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$). We adhered to the standard experimental setup found in the literature [3, 92, 4], employing a train/validation/test split of 200200/600600/$600$ for Walking and 200200/240240/$240$ for Running (additional details in Appendix C.3).
@@
We apply our training procedure REMUL to the Transformer architecture and compare it with SE(3)-Transformer, Equivariant Graph Neural Operator (EGNO) [4], Geometric Algebra Transformer (GATr), standard Transformer, and Transformer trained with data augmentation. We also compare with Equivariant MLP [93], as well as two approximate equivariance architectures: Residual Pathway Priors (RPP) [80], and Projection-Based Equivariance Regularizer (PER) [94]. As these architectures are designed specifically on MLP and linear layers, we apply our method to a standard MLP with a similar number of parameters. Our results are presented in Table 2. For REMUL, we provide plots on how the performance and equivariance error change w.r.t. the penalty parameter $\beta$ in Figure 2.
@@
Table 2: Motion Capture dataset: MSE ($\times 10^{-2}$). REMUL procedure and data augmentation (DA) were applied to standard Transformer and MLP. First, Second. REMUL comes best in both tasks.
[…]
@@
Table 2 indicates that when processing 3D positions related to human motions, both SE(3)-Transformer and GATr perform worse than the standard Transformer. This outcome is noteworthy because human motion often lacks full rotational symmetry, particularly along the vertical or gravity axis. In fact, as detailed in the Appendix D.5 (Table 9), our analysis of axis-specific equivariance errors for REMUL-Transformer confirms that the error is highest for rotations around the $Z$-axis. Consequently, imposing strict $SO(3)$ equivariance across all axes may not be beneficial and can be detrimental to performance. In contrast, a standard Transformer trained with REMUL has the best performance in both tasks. Following Figure 2, there is a noticeable trade-off: while higher $\beta$ values reduce overall equivariance error, optimal task performance is often observed at an intermediate level of learned equivariance, where the model balances between being too rigid (fully equivariant) and too flexible (non-equivariant). This underscores that the optimal degree of symmetry is task-dependent and that REMUL’s flexibility in learning approximate equivariance is advantageous for such real-world scenarios.
@@ 6.3 Molecular Dynamics
We also present a comparative analysis between constrained equivariant models and unconstrained models focusing on molecular dynamics, specifically predicting 3D molecule structures. We utilize the MD17 dataset [95], which comprises trajectories of eight small molecules. We use the same dataset split in [92, 4], allocating $500$ samples for train, $2000$ for validation, and $2000$ for test. For this task, we selected the Equivariant Graph Neural Network (EGNN) architecture and its non-equivariant GNN counterpart, as presented in [5]. We then apply REMUL procedure as well as data augmentation to the GNN architecture. Both architectures have the same hyperparameters (more information is indicated in Appendix C.4). We also compare with GMN [92], EGNO [4], and HEGNN [96]. Our results are provided in Table 3. We illustrate how the performance and equivariance error of a GNN trained with REMUL vary across different molecules as a function of $\beta$ in Figure 9 and Figure 10.
Table 3: MD17 dataset: MSE ($\times 10^{-2}$). REMUL procedure and data augmentation (DA) were applied to GNN. First, Second.
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic Toluene Uracil EGNN $14.41_{\pm 0.15}$ $62.40_{\pm 0.53}$ $4.64_{\pm 0.01}$ $13.64_{\pm 0.01}$ $0.47_{\pm 0.02}$ $02_{\pm 0.02} 11.$$1.78_{\pm 0.07}$ $0.64_{\pm 0.01}$ GMN $10.14_{\pm 0.03}$ $48.12_{\pm 0.4}$ $4.83_{\pm 0.01}$ $13.11_{\pm 0.03}$ $0.40_{\pm 0.01}$ $0.91_{\pm 0.01}$ $10.22_{\pm 0.08}$ $0.59_{\pm 0.01}$ EGNO $9.18_{\pm 0.06}$ $48.85_{\pm 0.55}$ $4.62_{\pm 0.01}$ $12.80_{\pm 0.02}$ $0.37_{\pm 0.01}$ $0.86_{\pm 0.02}$ $10.21_{\pm 0.05}$ $0.52_{\pm 0.02}$ HEGNN $9.94_{\pm 0.07}$ $59.93_{\pm 5.21}$ $4.62_{\pm 0.01}$ $12.85_{\pm 0.01}$ $0.37_{\pm 0.02}$ $0.88_{\pm 0.02}$ $10.56_{\pm 0.33}$ $0.54_{\pm 0.01}$ GNN $9.26_{\pm 0.40}$ $26.13_{\pm 0.11}$ $4.26_{\pm 0.03}$ $18.45_{\pm 0.54}$ $0.54_{\pm 0.001}$ $1.02_{\pm 0.02}$ $9.93_{\pm 0.82}$ $0.70_{\pm 0.001}$ DA-GNN $3.7_{\pm 0.04} 11$$10.93_{\pm 5.3}$ $5.74_{\pm 0.02}$ $13.65_{\pm 0.02}$ $0.69_{\pm 0.001}$ $1.33_{\pm 0.04}$ $19.14_{\pm 0.001}$ $0.73_{\pm 0.002}$ REMUL-GNN $9.28_{\pm 0.40}$ $25.95_{\pm 0.18}$ $4.02_{\pm 0.16}$ $13.59_{\pm 0.03}$ $0.54_{\pm 0.001}$ $0.99_{\pm 0.001}$ $9.38_{\pm 0.20}$ $0.67_{\pm 0.001}$
@@
From the results presented in Table 3, GNN trained with REMUL outperforms EGNN in six out of eight molecules. Interestingly, a standard GNN, without data augmentation or REMUL, surpasses the performance of EGNN on multiple molecules, such as Aspirin and Toluene. In Figure 9 & Figure 10, we observe that the optimal performance of each molecule is attained at different values of the penalty parameter $\beta$. For instance, Malonaldehyde exhibits a direct correlation between model performance and equivariance, where a higher $\beta$ yields better performance. Conversely, for most other molecules, there appears to be a pronounced trade-off where the best performance is achieved at a lower value of $\beta$. This is particularly evident with molecules like Aspirin, where a standard GNN architecture outperforms EGNN. We also plot the 3D structures of the eight molecules in Figure 12. Molecules such as Malonaldehyde, characterized by their symmetric components, might be ideally suited for equivariant design. However, this advantage does not apply to all molecules. Aspirin on the other side, might have an asymmetric structure and exhibit a range of interactions and dynamic states that equivariant models might simplify. Consequently, for such molecules, less equivariant models could potentially offer more accurate predictions.
Finally, while REMUL also achieves competitive performance relative to other equivariant models such as EGNO and HEGNN, it is important to note that these models incorporate architectural elements that enhance geometric properties in distinct ways and might not be directly comparable to a simple GNN. For example, EGNO employs additional Fourier features, while HEGNN leverages high-degree steerable features.